"A Watch t use 0.62/1.25 points 1 Prefous Anewe LaPCalc 10 B.018 Consider the follewing. x-3 cos 0 y5 sin 8 (a) Sketch the curve represented by the parametric equations (indicate the orlentation of the curve) y 10 -2 21 2 -10 21 -10 (b) Eliminate the parameter and write the corresponding rectangular equation whose graph represents the curve Adjust the domain of the resuting rectangular eqvation if necessary oed Heln?
0/1.25 points | Previous Answers LarPCalc8 10.6.042 6. Find a set of parametric equations for the rectangular equation using t x and t 4-x. x- 3y 2 (a) tx X y = (b) t 4-x X y= Need Help? Watch It Read It Talk to a Tutor Your work in question(s) 5 will also be submitted or saved Submit Aceinnment
O1.25 points LarPCalc8 10.6.044. 7. Find a set of parametric equations for the rectangular equation using t x and t 3-x. y x2+4 (a) tx X= y (b) t 3-x X y Need Help? Read It Watch It Talk to a Tutor A Your work in question(s) 5 will also be submitted or saved.
012 points LarPCalc8 10.5.014.MISA This question has several parts that must be completed sequentially. If you skip a part of the question, you will not receive any points for the skippe Tutorial Exercise Rotate the axes to eliminate the xy-term in the equation. Then write the equation in standard form. Sketch the graph of the resulting equation, showing both sets of axes. xy-6-0
2. O-2 points LarPCalc8 10.5.016 MLSA This question has several parts that must be completed sequentially, If you skip a part of the question, you will not receive any points for the skip Tutorial Exercise Rotate the axes to eliminate the xy-term in the equation. Then write the equation in standard form. Sketch the graph of the resulting equation, showing both sets of axes. xy + 2x y+5 0
Answer #1

WELCOME

GIVEN

X = 3cosθ

So cosθ = x/3

Y = 5sinθ

So sinθ = y/5

We know that

Sin2θ + cos2θ = 1

Or (x/3)2 + (y/5)2 = 1

Or x2 / 32 + y2 / 52 = 1

Thus   a = 3 and b = 5

So ‘a’ is the correct graph